OpenStudy (anonymous):

The formula for determining interest compounded monthly is A = P(1 +r/12 )^(12t), where A represents the amount invested after t years, P the principal invested, and r the interest rate. Jimmy invests $2,000 at an interest rate of 10% for 4 years, while Jenny invests $2,000 at an interest rate of 5% for 8 years. Determine the amount of return gained by Jimmy and Jenny. Summarize your results from Part 1, including how you arrived at your answer.

OpenStudy (kira_yamato):

For Jimmy: P = 2000, r = 0.1, t = 4 A(Jimmy) = 2000(1 + 0.1/12)^(12(4)) = $2978.71 Gain from interest = $(2978.71 - 2000) = $978.71 For Jenny: P = 2000, r = 0.05, t = 8 A(Jenny) = 2000 (1 + 0.05/12)^(12(8)) = $2981.17 Gain from interest = $(2981.17 - 2000) = $981.17